# RANA: Robust Active Learning for Noisy Network Alignment

### v1

@@ 3 Preliminaries
In network alignment, we focus on two networks: the source network $G^{s}=(V^{s},E^{s},\mathbf{A}^{s})$ with node attribute matrix $\mathbf{X}^{s}$, and the target network $G^{t}=(V^{t},E^{t},\mathbf{A}^{t})$ with node attribute matrix $\mathbf{X}^{t}$. $V^{s}$ and $V^{t}$ represent the node sets, and $E^{s}$ and $E^{t}$ represent the edge sets. $N^{s}=|V^{s}|$, $N^{t}=|V^{t}|$. The source and target networks are related by a pre-alignment matrix $\mathbf{P}\in\{0,1\}^{N^{s}\times N^{t}}$, where $[\mathbf{P}]_{ij}=1$ indicates that $v_{i}^{s}\in V^{s}$ and $v^{t}_{j}\in V^{t}$ is a pair of pre-aligned anchor nodes. In practice, the pre-alignment matrix is typically extremely sparse, providing only limited supervision, label sparsity emerges as a key challenge in network alignment. The network alignment model predicts anchor links, typically represented by an alignment matrix $\mathbf{S}$, where $[\mathbf{S}]_{uv}$ indicates the similarity between nodes $u\in V_{s}$ and $v\in V_{t}$. Formally, we define the network alignment problem as follows:
@@
Network Alignment: Given two networks $G^{s}$ and $G^{t}$ with a pre-alignment matrix $\mathbf{P}$, the goal of network alignment is to calculate an alignment matrix $\mathbf{S}$ where $[\mathbf{S}]_{ij}>[\mathbf{S}]_{ir|r\neq j}$ iff there exists an anchor link between node $v^{s}_{i}$ and $v_{j}^{t}$.
@@ 4.3 Noise-aware Selection Module
Influence Score: The influence score of node $v_{i}$ on node $v_{j}$ after $k$-step propagation is the L1-norm of the expected Jacobian matrix $\hat{I}(v_{j},v_{i},k)=||\mathbb{E}[\partial{\mathbf{X}_{j}^{(k)}}/\partial{\mathbf{X}_{i}^{(0)}}]||_{1}$. Formally, the normalized influence score is defined as
@@ 3 Train network alignment model with $\mathcal{V}_{train}$;
4 Get prediction probability $p_{ij}$, labels $\hat{y}_{ij}$ and Accuracy $Acc$;
@@ C.2 Baseline Details
[…]
@@
[…]
@@ C.3 Evaluate Metrics
[…]

### v2

@@ 3 Preliminaries
In network alignment, we focus on two networks: the source network $G^{s}=(V^{s},E^{s},\mathbf{A}^{s})$ with node attribute matrix $\mathbf{X}^{s}$, and the target network $G^{t}=(V^{t},E^{t},\mathbf{A}^{t})$ with node attribute matrix $\mathbf{X}^{t}$. $V^{s}$ and $V^{t}$ represent the node sets, and $E^{s}$ and $E^{t}$ represent the edge sets. $N^{s}=|V^{s}|$, $N^{t}=|V^{t}|$. The source and target networks are related by a pre-alignment matrix $\mathbf{P}\in\{0,1\}^{N^{s}\times N^{t}}$, where $\mathbf{P}]_{ij}=1$ indicates that $v_{i}^{s}\in V^{s}$ and $v^{t}_{j}\in V^{t}$ is a pair of pre-aligned anchor nodes. In practice, the pre-alignment matrix is typically extremely sparse, providing only limited supervision, label sparsity emerges as a key challenge in network alignment. The network alignment model predicts anchor links, typically represented by an alignment matrix $\mathbf{S}$, where $\mathbf{S}]_{uv}$ indicates the similarity between nodes $u\in V_{s}$ and $v\in V_{t}$. Formally, we define the network alignment problem as follows:
@@
Network Alignment: Given two networks $G^{s}$ and $G^{t}$ with a pre-alignment matrix $\mathbf{P}$, the goal of network alignment is to calculate an alignment matrix $\mathbf{S}$ where $\mathbf{S}]_{ij}>[\mathbf{S}]_{ir|r\neq j}$ iff there exists an anchor link between node $v^{s}_{i}$ and $v_{j}^{t}$.
@@ 4.3 Noise-aware Selection Module
[…]
@@ 3 Train network alignment model with $\mathcal{V}_{train}$;
4 Get prediction probability $p_{ij}$, labels $\hat{y}_{ij}$ and Accuracy AccAcc;
@@ C.2 Baseline Details
  • TopMatchings [20]: In the TopMatchings method, node determinism is quantified by constructing a weighted bipartite graph and calculating the top-$l$ maximum weight matching. The determinism of node $v$ is defined as $\text{Cert}(v)=\underset{u\in\mathcal{C}_{v}}{\max}P(M(v)=u|\mathcal{H})$, where $P(M(v)=u|\mathcal{H})$ represents the marginal distribution of node $v$ matching node $u$, based on the sampled matching set $\mathcal{H}$. If the distribution of node $v$ matching different nodes in the sample is fairly uniform, it indicates higher uncertainty in the best match for the node, i.e., lower determinism. TopMatchings selects the node with the lowest determinism for query, i.e., $\hat{v}=\underset{v\in\mathcal{V}_{s}}{\arg\min}\text{Cert}(v)$.
@@
  • Least Confident: An intuitive query strategy in active learning is to select the node with the least confidence in its prediction. We define the information of the query node as $I(v)=1-\underset{u}{\max}{\mathbf{X}(u,v)}$, and therefore select the node with the maximum information for query, i.e., $\hat{v}=\underset{v}{\max}I(v)$.
  • Margin: Nodes with the smallest difference between the maximum and minimum predicted values are considered to have the highest alignment uncertainty. We define the information of the query node as $I(v)=\mathbf{X}(m,v)-\mathbf{X}(n,v)$, where $\mathbf{X}(m,v)$ and $\mathbf{X}(n,v)$ are the maximum and minimum values of $\mathbf{X}$, respectively. We select the node with the smallest difference for query, i.e., $\hat{v}=\underset{v}{\min}I(v)$.
@@ C.3 Evaluate Metrics
We use Acc@k [40] and MAP [21] to evaluate the performance of our framework. Acc@k indicates if a node’s true anchor match is present in a list of top-$k$ potential anchors.